OpenStudy (brucebaner):

What is the value of expression? 6−7 • 64 A. 6−3 B. 6−11 C. 6−28 D. 6^11

OpenStudy (brucebaner):

@SolomonZelman

OpenStudy (brucebaner):

@ganeshie8

OpenStudy (brucebaner):

help me plz @wio

OpenStudy (brucebaner):

@AriPotta

OpenStudy (brucebaner):

help me @AriPotta and @wio

OpenStudy (aripotta):

is that the exact expression or are there any typos?

OpenStudy (brucebaner):

okay

OpenStudy (aripotta):

ok what?

OpenStudy (brucebaner):

so can you help me

OpenStudy (brucebaner):

yes

OpenStudy (aripotta):

so it's 6 minus 7 times 64?

OpenStudy (brucebaner):

so its C

OpenStudy (aripotta):

i'm asking if that's the correct problem, because i don't see a correct answer

OpenStudy (brucebaner):

oh okay its 6^-7*6^4

OpenStudy (aripotta):

ah yes. that helps :P \[6{^{-7}} =~\frac{ 1 }{ 6^{7} }\] so \[\frac{ 1 }{ 6^{7} }\times 6^{4}\]

OpenStudy (brucebaner):

okay its C

OpenStudy (aripotta):

what are the choices?

OpenStudy (brucebaner):

A. 6^−3 B. 6^−11 C. 6^−28 D. 6^11

OpenStudy (aripotta):

we're basically solving \[\frac{ 6^{4} }{ 6^{7} }\]

OpenStudy (brucebaner):

A

OpenStudy (aripotta):

yea, i'd say a too :)

OpenStudy (brucebaner):

lol

OpenStudy (brucebaner):

bye thx for helping me @AriPotta

OpenStudy (aripotta):

\(\Large \boldsymbol{ \rlap{\color{#008AFF}{\hspace{2.5pt}n}\color{#0073FF}{\hspace{0.34pt}o} \hspace{3pt}\color{#005CFF}{p} \color{#0045FF}{\hspace{0.9pt}r} \color{#0017FF}{o} \color{#0200FF}{\hspace{1.7pt}b} \color{#3000FF}{\hspace{1.5pt}l} \color{#4700FF}{\hspace{0.4pt}e} \color{#7500FF}{\hspace{0.9pt}m} \color{#8C00FF}{\hspace{0.5pt}o}~\LARGE\color{#D100FF}{\ddot\smile}}{\hspace{1.5pt} no~p\hspace{0.3pt}r\hspace{0.4pt}o\hspace{1pt}b\hspace{1.5pt}l\hspace{0.4pt}e\hspace{0.4pt}m\hspace{0.4pt}o}}\)